OpenStudy (anonymous):

lim (1 + 1/x)^(3x) as x->infinity solving this using L'hopital's rule for indeterminate infinity - infinity.

OpenStudy (anonymous):

lim ln([(x+1)/x]^3x) as x ->.infinity = lim ln([(x+1)^(3x)]/[x^(3x)]) as s->infinity = lim ln((x+1)^(3x))-ln(x^(3x)) = infinity - infinity is this a correct start?

OpenStudy (anonymous):

Don't worry this is only calc 2

OpenStudy (anonymous):

you should probably just recognize this number

OpenStudy (anonymous):

\[\lim_{x\to \infty}(1+\frac{1}{x})^x=e\]

OpenStudy (anonymous):

and so your answer is \(e^3\) but you can use l'hopital if you like take the log, get \[3x\ln(1+\frac{1}{x})\]which is in the form \(\infty\times 0\) then use the usual trick of rewriting as \[\frac{\ln(1+\frac{1}{x})}{\frac{1}{3x}}\]and use l'hopital for this one

OpenStudy (anonymous):

in answer to your question, no, that was not the correct start

OpenStudy (anonymous):

ah thank you

OpenStudy (anonymous):

yw